A Bibliography of Publications of Cleve B. Moler

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 April 2022 Version 1.59

Abstract

This bibliography records publications of Cleve B. Moler.

Title word cross-reference

10^{435} [Mol95b]. 2 [Mol74]. Ax = λBx [MS71]. AXB^{T} + CXD^{T} = E [GWL^{+}92, GLAM92].

-by- [Mol74].

1 [Mol18a]. 128-bit [Mol17c]. 15th [Mol92]. 16-bit [Mol17b]. 1978 [MS78a].

2007 [DFM^{+}07]. 2012 [Ano12].

50th [DFM^{+}07].

6 [Mol00a]. 6600 [DM73]. 6600/7600 [DM73].

754 [Mol19b]. 75th [DFM^{+}07]. 7600 [DM73]. 77 [GWL^{+}92].

A&M [Dut16]. Accuracy [DMW83]. Accurate [Mol67b]. ACM [ACM72]. acoustics [MS70b]. Affair [CMMP95]. Algebra [GM87a, HM88, Mol80d]. Algebraic [FM67, Joh68, FM71, Mol67b, MS78a]. algebraische [FM71]. Algorithm [GWL^{+}92, MS73, MS78b, MS71, MS72, Mol95c, Mol01]. algorithms [MS78a]. Analysis [GO80, MM83a, Mol68b, Mol68c, Mol69c, Mol69d, Mol83, Mol70]. Analyst [Hig99]. Analytic [LM67, MRS94]. anniversary [DFM^{+}07]. Annual [ACM72]. anymore [Mol15b]. apology [Mol99a]. Applications [GM87a]. approximate [Efi64]. Approximations [FHM67]. April [AFI67]. Ardent [Mol18a]. Arithmetic [Mol17b, Mol17c, Mol19b, Mol19c, Mol96a]. arrival [DFM^{+}07]. Art [DFM^{+}07, Mol69e]. Aspects [CMMP95]. Atlantic [AFI67]. August [ACM72, MS78a]. Award [Ano11].

B [AF69, Joh68]. basis [MRS94]. Bauer [MB96, Mol15a]. be [GM87b]. Beebe [Mol79]. Before [Mol19b]. Behavior [Mol01]. Benchmarks [Mol94]. Beresford [Mol08a]. billion [Mol02b]. birthday [DFM^{+}07, Mol92]. bit [Mol17b, Mol17c]. BLAS [Mol88b, Mol88c]. Book [AF69, CMR73, Cur78, Joh68]. Boston [ACM72]. boundary [Efi64]. Bounds [FHM67, MP68]. Broadwell [GM91]. Bug [Mol17a]. Building [DGG^{+}08]. Burton [Mol70].

C [AF69, Cur78]. calculations [Efi64]. calculator [Mol80a]. can [GM87b]. capabilities [Mol00b]. CDC [DM73]. celebrating [DFM^{+}07]. certification [Mol63]. certify [BCC^{+}72]. chance [Mol95a]. Chemistry [Mol79, MS78a]. City [AFI67]. Class [Dub83]. Cleve [Ano12, Joh68, Ano11, Dut16, Hai08, Hig99, Mar99, Mol94, Mol95c, Mol95b, Mol95a, Mol95d, Mol96a, Mol96b, Mol97, Mol98a, Mol98b, Mol99b, Mol99a, Mol00a, Mol00b, Mol01, Mol02a, Mol02b, Mol06b, Mol06a, Mol13, Mol14b, Mol14a, Mol15a, Mol17a, Mol17b, Mol17c, Mol18a, Mol19a, Mol19b, Mol19c]. Cleve s [Mol19d]. Cody [MM09]. collaborative [BCC^{+}72]. Collisionless [GM91]. Community [DGG^{+}08]. Compiling [JM94]. Complex [Mol98b]. comprehensive [Mol68a]. Computation [Dub83, Mol02b, LM91, Mol00b]. Computational [CMMP95, DFM^{+}07, HM88, Jam89, PAB^{+}01]. Computations [FMM77, FMM80, Mol72b, Mol86, Mol69e, Mol71, Ste68, Cur78]. Compute [MV78a, MV78b, MV03, MV79]. Computed [DMW83]. Computer [AFI67, Ano12, FM67, FMM77, MLB87c, FM71, FMM80, Cur78, Joh68]. Computer-Verfahren [FM71]. computers [DM73]. Computing [DGG^{+}08, DFM^{+}07, Mol19a, Mol75, Mol83, Mol04a, Mol04b, Mol08b]. Condition [CMSW79]. Conference [ACM72, AFI67, Hea86, DFM^{+}07]. Conjectured [CMR73]. Controversial [Mol14a]. Converging [Jam89]. Corner [Mol94, Mol95c, Mol95b, Mol95a, Mol95d, Mol96a, Mol96b, Mol97, Mol98a, Mol98b, Mol99b, Mol99a, Mol00a, Mol00b, Mol01, Mol02a, Mol02b, Mol06b, Mol06a, Mol13, Mol17a, Mol17b, Mol17c, Mol18a, Mol19a, Mol19b, Mol19c, Mol69a, Mol14b, Mol14a, Mol15a]. CR [SM22]. Craftsmanship [Mol02a]. Cramer [Mol74]. Creator [Dut16, Hai08]. crossing [Mol97]. Cryptograms [MM83a].

Decades [Mol06b]. Decomposing [MS78b]. Dedication [GDM87, Mol92]. Definite [CMR73]. Demonstration [Mol82b]. Denormals [Mol14a]. Design [GMS92, Mol80a]. Development [Cow84]. Difference [Mol65]. differential [Efi64, Mol96b, Mol97]. Differentiation [LM67]. Difficulties [Mol75]. digital [MS79]. Directions [DFM^{+}07]. disseminate [BCC^{+}72]. Distributed [Mol86, PPM^{+}89, Mol01]. draw [PPM^{+}89]. Dubious [MV78a, MV78b, MV03, MV79].

education [PAB^{+}01]. Efficiencies [Jam89]. efficient [MS79]. effort [BCC^{+}72]. Eigensystem [GBDM77, SBD^{+}76]. Eigenvalue [DM84, MS73, Mol79, MS71, MS72]. Eigenvalues [DMW83, FHM67, Mol65, MP68, Nic67, DM78, Efi64, Mol98a]. eigenvector [Mol02b]. Eigenvectors [DMW83, MP68, Efi64]. eigshow [Mol98a]. EISPACK [DM84, GBDM77, SBD^{+}76]. Elimination [SM22]. Elliptic [FHM67, Nic67]. engineering [PAB^{+}01]. Equation [GWL^{+}92, GLAM92, Mol72a, Mol96b]. equations [Efi64, GM91, Mol97]. era [Mol02a]. error [DM73, Mol95a]. Estimate [CMSW79]. event [Mol97]. examination [Mol68a, Mol68b, Mol68c, Mol69c, Mol69d]. Excerpts [Mol68a]. Experiments [HM88, Mol11]. explain [Mol98a]. Exponential [MV78a, MV78b, MV03, Mol75, MV79]. Extension [GBDM77, Nic67].

factorization [MS79]. Fame [Mol94]. Faster [Mol00a]. Fe [USE94]. features [Mol99b]. Fernbach [Ano11]. FFTW [Mol00a]. Finding [Jam89]. Finite [Mol65, Mol00a]. First [Mol70]. Fitting [Hig99]. Five [MV03]. Floating [Mol67a, Mol14b, Mol14a, Mol17a, Mol17b, Mol17c, Mol19b, Mol19c, Mol98b, Mol96a]. floating-point [Mol98b]. Format [Mol19c]. Formulae [Jam89]. Forsythe [Cur78, DFM^{+}07, Joh68, Mol72c]. FORTRAN [GWL^{+}92, Mol71, Mol72b]. FORTRAN-77 [GWL^{+}92]. fortune [Mol94]. four [GM91]. Fourier [Mol00a]. Fox [MS78b, Nic67]. Friedrich [Mol15a]. Function [Mol02a]. Functions [LM67, Mol98b, Mol02a]. Future [DFM^{+}07].

G [Cur78]. Gatlinburg [MB96, Mol13]. Gene [DFM^{+}07, Mol92]. Generalized [MS73, Mol79, MS71, MS72]. generates [Mol01]. generator [MM73]. geometrical [MS70b]. George [DFM^{+}07, Joh68, Mol72c]. German [FM71]. Golden [Mol96b]. Golub [DFM^{+}07, Mol92]. Good [Mol98a]. Google [Mol02b]. got [Mol88b, Mol88c]. Graduate [PAB^{+}01]. graphics [PPM^{+}89]. groupie [Mol07]. Growth [Mol06b]. Guide [DBMS79, GBDM77, Mol80c, Mol81, Mol82a, MLB87b, MLB87a, SBD^{+}76, CM93, LM91, ML90, Mol18b].

H [GM87a, Mol92]. Half

[Mol17a, Mol17b, Mol19c]. Half-Precision [Mol17a]. Hall [Cur78]. handling [Mol97]. Harmful [GM87b]. helps [Mol98a]. Henrici [Nic67]. High [USE94, LM91]. high-performance [LM91]. Historic [Mol18b]. history [ML20]. Householder [MS78b, MB96, Mol13]. hyperbolic [Mol98b]. Hypercube [Hea86].

IEEE [Ano12, Mol96a, Mol19b]. image [MS79]. Implementation [GMS92]. Improving [DMW83]. Incorporates [Mol00b, Mol00a]. Increasing [Mol00b]. Inequality [AF69]. infallibility [Mol95c]. Insignificant [Mol14a]. Installation [Mol81]. Integrating [MS70a]. integration [MS70b]. Interactive [Mol80b, Mol80a, Mol83]. Interview [Mar99]. inverse [Mol98b]. inversion [Mol63]. isn’t [Mol95d]. Iterative [Jam89, Mol67a].

J [AF69]. James [GM87a]. Jim [MM09]. Joint [AFI67]. just [Mol15b].

Knoxville [Hea86].

Laboratory

[Mol80b, Mol82b, Mol19d, Mol88a, Mol15b]. Languages [Edw09, USE94]. LAPACK [Mol00b]. Laplace [Mol65]. large [Mol95a, Mol99b]. large-scale [Mol99b]. Largest [Mol02b]. Later [MV03]. Learning [Mol83]. Level [USE94]. libraries [MM09]. library [GMI85]. Linear [FM67, GM87a, Joh68, Mol72a, Mol80d, FM71, Mol67b]. lineare [FM71]. lines [PPM^{+}89]. LINPACK [DBMS79, Mol94]. LOBO.LIB [GMI85]. long [Mol95b]. LU [SM22].

Lyness [AF69].

M [Cur78]. making [Mol95a]. Malcolm [Cur78]. March [DFM^{+}07]. Masinnye [FMM80]. matematiceskich [FMM80]. Math [Mar99, CM93]. Mathematical [Cow84, Cur78, FMM77, FMM80, Hai08, BCC^{+}72, GMI85, Mol88a, Mol02a, MM09]. Mathematics [DFM^{+}07, Mol19a, Mol83]. MathWorks [MRS94, Mol06b]. MATLAB [Dut16, Edw09, GMS92, LM91, Mol80b, Mol80c, Mol81, Mol82a, MLB87c, Mol88a, ML90, Mol94, Mol95d, Mol96b, Mol00a, Mol00b, Mol02a, Mol04a, Mol04b, Mol06b, Mol08b, Mol11, Mol18b, ML20, Hai08, JM94, MLB87b, MLB87a]. Matrices [GMS92]. Matrix [CMSW79, DM84, GBDM77, GWL^{+}92, GLAM92, HM88, Mol71, Mol72b, MS73, MV78a, MV78b, Mol80b, Mol86, Mol02b, MV03, SBD^{+}76, DM78, Mol63, Mol69b, Mol69e, MS71, MS72, Mol75, MS79, MV79, Mol80a, Mol82b, Mol00b, Mol15b, CMR73]. Meetings [MB96]. megaflops [Mol94]. Membrane [Mol18a]. Memory [GM87a, Mol86, Mol72c]. Methods [Cur78, Dub83, FMM77, FMM80, KMN89, Mol65, MS78a, Mol99b]. metody [FMM80]. Mexico [USE94]. model [Mol96a]. Moler [AF69, Cur78, Joh68, Ano11, Ano12, Dut16, Hai08, Hig99, Mar99, Mol79, Mol19a, Nic67]. Multiprocessors [Mol86, Hea86].

N [AF69]. NA-Net [DGG^{+}08]. NATS [BCC^{+}72]. Net [DGG^{+}08]. Netlib [DGG^{+}08]. Nineteen [MV78b, MV79, MV03]. NJ [AFI67]. Normal [Mol01]. normally [Mol01]. Notes [Jam89]. Number [CMSW79, MM73]. Numbers [Mol95a, Mol14b, Mol95e, Mol01]. numeric [LM91]. Numerical [DFM^{+}07, Dub83, GO80, Hig99, KMN89, LM67, Mol69b, Mol70, Mol80d, Mol02a, Mol04a, Mol08b, Mol68b, Mol68c, Mol69c, Mol69d, MS70b, Mol83, MRS94, Mol04b, MS78a].

Object [Mol17a]. Objectively [Mol99a]. occasion [Mol92]. October [Hea86, USE94]. ODEs [Mol96b]. OOPS [Mol99a]. Operator [Mol65]. Operators [FHM67, MP68, Nic67]. Optimally [Mol99b]. Optimization [Mol99b]. order [Mol02b]. ordinary [Mol96b].

Package [DM84, GWL^{+}92]. PageRank [Mol02b]. pages [Cur78]. Paging [Mol72b, Mol71]. Paper [Nic67]. parallel [Mol95d]. partial [Efi64]. PC [MLB87b]. PC-Matlab [MLB87b]. Pentium [CMMP95, Mol95a]. performance [LM91]. Perspective [Mol15b]. Ph.D. [Mol68a, Mol68b, Mol69c]. Pioneer [Ano12, Hai08, MM09]. Pioneering [Hig99]. Point [Mol67a, Mol14b, Mol14a, Mol17a, Mol17b, Mol17c, Mol19b, Mol19c, Mol98b]. points [Mol96a]. Positive [CMR73]. Precision [Mol17a, Mol17b, Mol17c, Mol19c]. Prentice [Cur78]. Prentice/Hall [Cur78]. Principles [Mol70]. PRO [LM91, MLB87c, ML90, MLB87a]. PRO-MATLAB [LM91, MLB87c, ML90, MLB87a]. Problem [Kah71, Mol79, MS71]. Problems [DM84, ML67, MS73, Mol80d, Efi64, Mol69b, MS72, Mol99b]. procedures [Mol63]. Proceedings [ACM72, Hea86, USE94]. processing [MS79]. Professor [Mol06a]. Programming [Edw09]. progress [Mol04b]. Projection [MS78b]. Purdue [Mol68c, Mol69d]. Putting [Mar99]. Pythagorean [Dub83, MM83b].

QR [Mol95c]. Quadruple [Mol17c].

qualifying

[Mol68b, Mol68c, Mol69c, Mol69d].

Quantum [Mol79].

Random [Mol95b, MM73, Mol01]. Rapidly [Jam89]. Recollections [Mol07]. Refinement [Mol67a]. rely [Mol02a]. Remark [Mol63]. Remembering [Mol08a]. Remembers [MB96]. Replacing [MM83b]. report [MS78a]. Research [Mol80d]. Response [Mol79]. Restores [Mol18a]. Review [AF69, CMR73, Joh68, Cur78]. Reviews [Mol70]. Revisiting [Mol98b]. rid [Mol88b, Mol88c]. Roots [Jam89, MM83b, MS70a]. Roundoff [DM73]. Routines [GBDM77, SBD^{+}76]. rule [Mol74]. Russian [FMM80].

s [Mol19a]. Santa [USE94]. scale [Mol99b]. Schmidt [CMR73]. science [PAB^{+}01]. Scientific [DGG^{+}08]. Second [Hea86]. Sensitivity [DM78]. September [Hea86]. Short [Jam89]. simulations [MRS94]. SIMULINK [Mol96b]. Single [Mol88b, Mol88c]. Single-user [Mol88b, Mol88c]. Singular [MM83a, Mol98a]. skills [Mol02a]. small [Mol95a]. Software [Cow84, GWL^{+}92, Hai08, KMN89, BCC^{+}72, GMI85, LM91, MM09]. Solution [FM67, GLAM92, Joh68, FM71, Mol67b, Mol69b]. solutions [GM91]. Solver [Mol72a]. solvers [Mol96b]. Solving [DM84, GWL^{+}92]. Sources [Cow84]. Sparse [GMS92]. Speaking [Mol99b, Mol99a]. special [Mol02a]. speed [Mol00b]. sphere [Mol69a]. splines [MS70b]. Spring [AFI67]. Square [Jam89, MM83b, MS70a]. Standard [Mol96a]. Stanford [DFM^{+}07, DFM^{+}07, Mol07]. State [DFM^{+}07, Mol69e]. Striving [Mol95c]. Subject [Mol98b]. Sums [Dub83, MM83b]. Sun [ML90]. supercomputers [Mol88b, Mol88c]. survey [Mol67b]. SVD [Mol06a]. Sylvester [GWL^{+}92, GLAM92]. Symbolic [CM93]. Symmetric [MP68]. Symposia [Mol13]. Symposium [USE94]. Systeme [FM71]. Systems [FM67, FM71, Mol67b, Mol74, Joh68].

Tale [Mol95a, Mol95e]. teaching [Mol04b]. Technical [Mol15b]. Tennessee [Hea86]. Tetragamma [Mol02a]. Texas [Dut16]. their [Jam89]. there [Mol95a, Mol95d, Mol97]. thoughts [Mol95b]. Three [Mol80d]. time [Mol95b]. Titan [Mol18a]. toolbox [CM93, Mol99b]. tools [MRS94]. Transforms [Mol00a]. trap [Ste68]. Tribute [Hig99]. trig [Mol98b]. Trigonometric [AF69]. Trigonometry [Mol98b]. Twenty [MV03]. Twenty-Five [MV03]. Two [Mol95a, Mol06b, Mol95e].

Underflow [GM87b]. unifies [Mol96a].

universal [MM73]. University [DFM^{+}07, Dut16, Mol68c, Mol69d]. URAND [MM73]. USA [USE94]. Use [MS70b]. User [DBMS79, Mol80c, Mol82a, MLB87a, Mol18b, CM93, LM91, Mol88b, Mol88c, ML90]. Users [MLB87b].

Value [MM83a, Efi64]. values [Mol98a]. Variable [Mol19c]. VAX [LM91, MLB87c]. VAX/VMS [MLB87c]. VAXBARN [Mol18a]. velocity [GM91]. Velvel [Mol08a]. Verfahren [FM71]. version [MLB87c]. Very [USE94, Mol95b, Mol95a]. VHLL [USE94]. Vibrating [Mol18a]. Vibrations [Mol98a]. Visits [Dut16]. visualization [Mol88a]. VMS [MLB87c]. vycislenij [FMM80].

W [CMR73]. Ways

[MV78a, MV78b, MV03, MV79]. Wendroff [Mol70]. where [PPM^{+}89]. Wilkinson [GM87a]. Wins [Ano11]. Work [Mar99, Mol04b]. workshop [MS78a]. workstations [ML90]. World [Mol02b].

Years [MV03, Mol95b].

Z [Edw09]. Zero [Mol97]. Ziggurat [Mol01].

References

ACM:1972:PAA

[ACM72] Proceedings of the ACM Annual Conference, August 1972, Boston. ACM Press, New York, NY 10036, USA, 1972. LCCN TK 7885 A84p 1972. Two volumes.

Askey:1969:BRB

[AF69] Richard Askey and James Fitch.

Book review: A Trigonometric Inequality (J. N. Lyness and C. B. Moler). SIAM Review, 11 (1):82–86, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

AFIPS:1967:ACP

[AFI67] 1967 Spring Joint Computer Conference, April 18–20, Atlantic City, NJ, volume 30 of AFIPS conference proceedings. Thompson Book Co., Washington, DC, USA, 1967. LCCN TK7885.A1 J6 1967.

Anonymous:2011:CMW

[Ano11] Anonymous. Cleve Moler wins Fernbach Award. Computer, 44 (11):71, November 2011. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2012:ICP

[Ano12] Anonymous. 2012 IEEE Computer Pioneer to Cleve Moler. IEEE Web site., 2012. URL

https://www.computer.org/profiles/

cleve-moler. For improving the quality of mathematical software, making it more accessible and creating MATLAB.

Boyle:1972:NCE

[BCC^{+}72] J. M. Boyle, W. J. Cody, W. R. Cowell, B. S. Garbow, Y. Ikebe, C. B. Moler, and B. T. Smith. NATS, a collaborative effort to certify and disseminate mathematical software. In ACM ’72 [ACM72], pages 630–635. LCCN TK 7885 A84p 1972. Two volumes.

Chen:1993:SMT

[CM93] Denise Chen and Cleve B. Moler. Symbolic math toolbox: user’s guide. MathWorks, Natick, MA, USA, August 1993. (various) pp. LCCN QA76.73.M38 C454 1993.

Coe:1995:CAP

[CMMP95] Tim Coe, Terje Mathisen, Cleve Moler, and Vaughan Pratt. Computational aspects of the Pentium affair. IEEE Computational Science & Engineering, 2(1):18–30, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=372929;
 http://www.computer.org/cse/
cs1998/c1018abs.htm.

Ciment:1973:BRB

[CMR73] Melvyn Ciment, Cleve Moler, and Otto G. Ruehr. Book review: A Conjectured Positive Definite Matrix (K. W. Schmidt). SIAM Review, 15(1):222–223, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cline:1979:ECN

[CMSW79] A. K. Cline, C. B. Moler, G. W.

Stewart, and J. H. Wilkinson. An estimate for the condition number of a matrix. SIAM Journal on Numerical Analysis, 16(2): 368–375, April 1979. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cowell:1984:SDM

[Cow84] W. Cowell, editor. Sources and Development of Mathematical Software. Prentice-Hall series in computational mathematics, Cleve B. Moler, Advisor. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1984. ISBN 0-13-823501-5. xii + 404 pp. LCCN QA76.95 .S68 1984. US$32.50.

Curran:1978:BRB

[Cur78] M. P. J. Curran. Book review: Computer Methods for Mathematical Computations, by G. Forsythe, M. Malcolm, and C. Moler, 1977; 259 pages. (Prentice/Hall, £12.80). The Computer Journal, 21(4):336, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
21/4/336.2.full.pdf+html.

Dongarra:1979:LUG

[DBMS79] J. J. Dongarra, J. R. Bunch, C. B.

Moler, and G. W. Stewart. LIN-PACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1979. ISBN 0-89871-172-X (paperback). 320 pp. LCCN QA76.73 .L22 L5 1979; QA184 .L56 1982; QA214 .L56 1979.

Du:2007:SSA

[DFM^{+}07] Ding-Zhu Du, Charles Farbat, Walter Murray, Michael Overton, Haesun Park, Michael Saunders, and James Varah, editors. STANFORD 50: State of the Art & Future Directions of Computational Mathematics & Numerical Computing: A conference celebrating the 50th anniversary George Forsythe’s arrival at Stanford and Gene Golub’s 75th birthday, Stanford University, March 29–31, 2007. ????, ????, 2007. ISBN ???? LCCN ????

Dongarra:2008:NNB

[DGG^{+}08] Jack Dongarra, Gene H. Golub, Eric Grosse, Cleve Moler, and Keith Moore. Netlib and NA-Net: Building a scientific computing community. IEEE Annals of the History of Computing, 30(2):30–41, April/June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4544554.

Dorr:1973:REC

[DM73] Fred W. Dorr and Cleve B. Moler. Roundoff error on the CDC 6600/7600 computers. ACM SIGNUM Newsletter, 8(2):24–26, April 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [Kah71] for the original problem.

Davis:1978:SME

[DM78] G. J. Davis and C. B. Moler. Sensitivity of matrix eigenvalues. International Journal for Numerical Methods in Engineering, 12 (9):1367–1373, ???? 1978. CO-DEN IJNMBH. ISSN 0029-5981 (print), 1097-0207 (electronic).

Dongarra:1984:EPS

[DM84] Jack J. Dongarra and Cleve B.

Moler. EISPACK: A package for solving matrix eigenvalue problems. In Cowell [Cow84], chapter 4, pages 68–87. ISBN 0-13-823501-5. LCCN QA76.95 .S68 1984. US$32.50.

Dongarra:1983:IAC

[DMW83] J. J. Dongarra, C. B. Moler, and J. H. Wilkinson. Improving the accuracy of computed eigenvalues and eigenvectors. SIAM Journal on Numerical Analysis, 20(1):23– 45, February 1983. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dubrulle:1983:CNM

[Dub83] Augustin A. Dubrulle. Class of numerical methods for the computation of Pythagorean sums. IBM Journal of Research and Development, 27(6):582–589, November 1983. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Dutta:2016:MCC

[Dut16] Sourav Dutta. MATLAB creator Cleve Moler visits Texas A&M University. SIAM News, 49(6):??, July/August 2016.

URL  https://sinews.siam.
org/DetailsPage/TabId/900/
ArtMID/2243/ArticleID/783/
MATLAB-Creator-Cleve-Moler-
Visits-Texas-AM-University.
aspx.

Edwards:2009:ZPL

[Edw09] Kathryn Edwards. The A–Z of programming languages: MAT-LAB. ComputerWorld, December 9, 2009. CODEN CMPWAB. ISSN 0010-4841. URL http:

//www.computerworld.com.au/
article/329191/a-z_programming_
languages_matlab/.

Efimenko:1964:ACE

[Efi64] V. A. Efimenko. On approximate calculations of the eigenvalues and eigenvectors of boundary value problems in partial differential equations. Technical report, Computer Science Department, Stanford University, Stanford, CA, USA, 1964. 20 pp. Translated by George Reiter and Cleve Moler from the original Russian, published in Izv. Akad. Nauk SSSR, Ser. Mat. 613–623 (1936).

Fox:1967:ABE

[FHM67] L. Fox, P. Henrici, and C. Moler. Approximations and bounds for eigenvalues of elliptic operators. SIAM Journal on Numerical Analysis, 4(1):89–102, March 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Forsythe:1967:CSL

[FM67] George E. Forsythe and Cleve B.

Moler. Computer Solution of Linear Algebraic Systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1967. xi + 148 pp. LCCN QA297 .F57 1967.

Forsythe:1971:CVL

[FM71] George E. (George Elmer) Forsythe and Cleve B. Moler. Computer-Verfahren für lineare algebraische Systeme. (German) [Computer solution of linear algebraic systems]. Verfahren der Datenverarbeitung. R. Oldenbourg, München, Germany, 1971. ISBN 3-486-33601-0. 168 pp. LCCN ???? Translated from the American original by Christine und Helmuth Späth.

Forsythe:1977:CMM

[FMM77] George E. (George Elmer) Forsythe, Michael A. Malcolm, and Cleve B. Moler. Computer Methods for Mathematical Computations. Prentice-Hall series in automatic computation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1977. ISBN 0-13-165332-6. xi + 259 pp. LCCN QA297 .F5681. US$16.95. Cited in Åke Björck’s bibliography on least squares, which is available by anonymous

ftp from math.liu.se in pub/
references.

Forsythe:1980:MMM

[FMM80] George E. Forsythe, Michael A.

Malcolm, and Cleve B. Moler. Mašinnye metody matematičeskich vyčislenij. (Russian) [Computer Methods for Mathematical Computations]. Izdatel’stvo Mir, Moscow, USSR, 1980. 280 pp. Translated from the English by Kh. D. Ikramov.

Garbow:1977:MER

[GBDM77] B. S. Garbow, J. M. Boyle, J. J. Dongarra, and C. B. Moler. Matrix Eigensystem Routines— EISPACK Guide Extension, volume 51 of Lecture Notes in Computer Science, Editors: G. Goos and J. Hartmanis. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-08254-9, 3-540-08254-9. viii + 343 pp. LCCN QA193 .M381, QA267.A1,L43 no. 51.

Golub:1987:D

[GDM87] Gene H. Golub, Iain Duff, and Cleve Moler. Dedication. Linear Algebra and its Applications, 88/89:1–12, 1987. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gardiner:1992:SSM

[GLAM92] Judith D. Gardiner, Alan J. Laub, James J. Amato, and Cleve B. Moler. Solution of the Sylvester matrix equation AXB^{T} + CXD^{T} = E. ACM Transactions on Mathematical Software, 18(2):223–231, June 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Golub:1987:LAA

[GM87a] G. H. Golub and C. Moler. Linear algebra and its applications — in memory of James H. Wilkinson. Linear Algebra and its Applications, 88–89:1–3, April 1987. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grosse:1987:UCB

[GM87b] Eric Grosse and Cleve Moler. Underflow can be harmful. SIAM News, 20(6):1, 1987. ISSN 0036-1437.

Greenberg:1991:CSF

[GM91] J. M. Greenberg and Cleve Moler.

Collisionless solutions to the four velocity Broadwell equations. In Multidimensional hyperbolic problems and computations (Minneapolis, MN, 1989), volume 29 of IMA Vol. Math. Appl., pages 140– 155. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.

George:1985:LLM

[GMI85] Karen George, Cleve B. Moler, and C. Iuzzolino. LOBO.LIB mathematical software library. Technical report CS85-1, Department of Computer Science, University of New Mexico, Albuquerque, NM, USA, January 23, 1985. iv + 50 pp.

Gilbert:1992:SMM

[GMS92] John R. Gilbert, Cleve Moler, and Robert Schreiber. Sparse matrices in MATLAB: Design and implementation. SIAM Journal on Matrix Analysis and Applications, 13 (1):333–356, January 1992. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Cited in Åke Björck’s bibliography on least squares, which is available by anonymous ftp from

                                        math.liu.se in pub/references.

Golub:1980:NA

[GO80] Gene H. Golub and Joseph Oliger, editors. Numerical Analysis, volume 22 of Proceedings of symposia in applied mathematics. American Mathematical Society, Providence, RI, USA, 1980. ISBN 0-8218-0122-8, 0-8218-9237-1 (ebook). viii + 135 pp. LCCN QA1 .S93 v.22.

Gardiner:1992:AFS

[GWL^{+}92] Judith D. Gardiner, Matthew R. Wette, Alan J. Laub, James J. Amato, and Cleve B. Moler. Algorithm 705: A FORTRAN-77 software package for solving the Sylvester matrix equation AXB^{T} + CXD^{T} = E. ACM Transactions on Mathematical Software, 18(2):232–238, June 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Haigh:2008:CMM

[Hai08] Thomas Haigh. Cleve Moler: Mathematical software pioneer and creator of Matlab. IEEE Annals of the History of Computing, 30(1):87–91, January/ March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Heath:1986:HM

[Hea86] M. T. Heath, editor. Hypercube multiprocessors, 1987: Proceedings of the Second Conference on Hypercube Multiprocessors, Knoxville, Tennessee, September 29–October 1, 1986. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1986. ISBN 0-89871-215-7. LCCN QA76.5 .C61921 1986.

Higham:1999:FTP

[Hig99] Nicholas J. Higham. A fitting tribute to a pioneering numerical analyst [Cleve Moler]. SIAM News, 32(10):3, December 1999. ISSN 0036-1437. URL https:/

/www.siam.org/news/news.php?

id=796. Celebrating Cleve Moler’s 60th birthday!

Hill:1988:ECM

[HM88] David R. Hill and Cleve B. Moler (consulting editor). Experiments in Computational Matrix Algebra. Random House, New York, NY, USA, 1988. ISBN 0-394-35678-0 (paperback). xii + 446 pp. LCCN QA188 .H55 1987.

Jamieson:1989:SNR

[Jam89] M. J. Jamieson. Short notes:

Rapidly converging iterative formulae for finding square roots and their computational efficiencies. The Computer Journal, 32 (1):93–94, February 1989. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
32/1/93.full.pdf+html;   http:
//www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_
01/tiff/93.tif;   http://www3.

                                        oup.co.uk/computer_journal/
                                        hdb/Volume_32/Issue_01/tiff/
                                        94.tif.

Johnson:1994:CM

[JM94] Stephen C. Johnson and Cleve Moler. Compiling Matlab. In USENIX Association [USE94], pages 119–127. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

Johnston:1968:BRB

[Joh68]  R. L. Johnston.  Book re-

view: Computer Solution of Linear Algebraic Systems (George E. Forsythe and Cleve B. Moler). SIAM Review, 10(3):384–385, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kahan:1971:P

[Kah71] W. Kahan. A problem. ACM SIGNUM Newsletter, 6(3):6, November 1971. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See response [DM73].

Kahaner:1989:NMS

[KMN89] David Kahaner, Cleve Moler, and Stephen Nash. Numerical Methods and Software. Prentice-Hall series in computational mathematics. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1989. ISBN 0-13-627258-4. xii + 495 pp. LCCN TA345 .K341 1989. US$50.

Lyness:1967:NDA

[LM67] J. N. Lyness and C. B. Moler. Numerical differentiation of analytic functions. SIAM Journal on Numerical Analysis, 4(2):202– 210, June 1967. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Little:1991:MHP

[LM91] John Little and Cleve B. Moler.

MATLAB high-performance numeric computation software: PRO-MATLAB user’s guide: PRO-MATLAB for VAX. Math-Works, South Natick, MA, USA, March 1, 1991. ???? pp. LCCN QA184.M3 L57 1991.

Marchioro:1999:ICM

[Mar99] Tom Marchioro. Interview: With Cleve Moler putting math to work. Computing in Science and Engineering, 1(4):10–??, July/ August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

/dlib.computer.org/cs/books/
cs1999/pdf/c4010.pdf.

Moler:1996:BRH

[MB96] Cleve Moler and Friedrich L.

Bauer. Bauer remembers Householder and the Gatlinburg meetings. NA Digest, 96 (27), July 18, 1996. URL

http://www.maths.manchester.
ac.uk/~higham/conferences/
householder/h96/bauertalk.
pdf;   http://www.netlib.org/
na-digest-html/96/v96n27.html;
 http://www3.math.tu-berlin.
de/householder_2008/Cleve.

htm. The article contains notes for F. L. Bauer’s after-banquet talk entitled Memories of Alston Householder (1904–1993), Householder Symposium XIII, June 17– 21, 1996 Pontresina, Switzerland.

Moler:1967:P

[ML67] C. B. Moler and J. N. Lyness. Problems. ACM SIGNUM Newsletter, 1(3):12–13, February 1967. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1990:PMS

[ML90] Cleve B. Moler and John Little, editors. PRO-MATLAB for Sun workstations: Matlab for Sun workstations: user’s guide. Math-Works, South Natick, MA, January 1990. (various) pp. LCCN TA345.5.P7 P75 1991; QA188 .P76 1990.

Moler:2020:HM

[ML20] Cleve Moler and Jack Little. A history of MATLAB. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):81:1–81:67, June 2020.

                                        URL https://dl.acm.org/doi/
                                        abs/10.1145/3386331.

Moler:1987:PMUb

[MLB87a] C. Moler, J. Little, and S. Bangert. Pro-Matlab User’s Guide. The MathWorks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1987. ISBN ???? LCCN ????

Moler:1987:PMUa

[MLB87b] C. B. Moler, J. N. Little, and S. Bangert. PC-Matlab Users Guide. Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1987.

Moler:1987:PMV

[MLB87c] Cleve Moler, John Little, and Steve Bangert. PRO-MATLAB for VAX/VMS computer, version 3.1. The MathWorks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1987. ISBN ???? LCCN ????

Malcolm:1973:UUR

[MM73] Michael A. Malcolm and Cleve B.

Moler. URAND: a universal random number generator. Technical Report STAN-CS-73-334, Computer Science Department, Stanford University, Stanford, CA, USA, 1973. 5 pp.

Moler:1983:SVA

[MM83a] C. B. Moler and D. Morrison.

Singular value analysis of cryptograms. American Mathematical Monthly, 90(2):78–87, 1983. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Moler:1983:RSR

[MM83b] Cleve Moler and Donald Morrison. Replacing square roots by Pythagorean sums. IBM Journal of Research and Development, 27(6):577–581, November 1983. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5390405;
 http://www.research.ibm.com/
journal/rd/276/ibmrd2706P.
pdf.

More:2009:JCP

[MM09] Jorge Moré and Cleve Moler. Jim Cody: Pioneer in mathematical software libraries. NA Digest postings, June 28, 2009.

                                        URL http://www.netlib.org/
                                        na-digest-html/09/v09n26.html#
                                        1.

Moler:1963:RCM

[Mol63] Cleve Moler. Remark on certification of matrix inversion procedures. Communications of the Association for Computing Machinery, 6(7):389, July 1963. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moler:1965:FDM

[Mol65] Cleve Barry Moler. Finite Difference Methods for the Eigenvalues of Laplace’s Operator. Ph.D. thesis, Department of Computer Science, Stanford University, Stanford, CA, USA, May 27, 1965. iv + 142 pp. URL http://

                                        search.proquest.com/docview/
                                        302329013?.

Moler:1967:IRF

[Mol67a] C. B. Moler. Iterative refinement in floating point. Journal of the Association for Computing Machinery, 14(2):316–321, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Moler:1967:ACL

[Mol67b] Cleve B. Moler. Accurate solution of linear algebraic systems— a survey. In AFIPS SJCC ’67 [AFI67], pages 321–324. LCCN TK7885.A1 J6 1967.

Moler:1968:EPD

[Mol68a] C. B. Moler. Excerpts from Ph.D. comprehensive examination. ACM SIGNUM Newsletter, 3(3):??, July 1968. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1968:PDQ

[Mol68b] Cleve Moler. Ph.D. qualifying examination: numerical analysis. ACM SIGNUM Newsletter, 3(1):??, April 1968. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1968:PUN

[Mol68c] Cleve Moler. Purdue University: numerical analysis qualifying examination. ACM SIGNUM Newsletter, 3(2):??, July 1968. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1969:MSC

[Mol69a] C. B. Moler. More on the sphere in the corner. ACM SIGNUM Newsletter, 4(1):7, January 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [Ste68].

Moler:1969:NSM

[Mol69b] C. B. Moler. Numerical solution of matrix problems. In The digest record of the 1969 joint conference on mathematical and computer aids to design, pages 15–26 (of xxv + 441). ????, ????, 1969.

Moler:1969:PDQ

[Mol69c] C. B. Moler. Ph.D. qualifying examination: numerical analysis. ACM SIGNUM Newsletter, 4(2):13–16, June 1969. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1969:PUN

[Mol69d] C. B. Moler. Purdue University: numerical analysis qualifying examination. ACM SIGNUM Newsletter, 4(1):11–13, January 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1969:SAM

[Mol69e] Cleve B. Moler. State of the art in matrix computations. ACM SIGNUM Newsletter, 4(1):22– 28, January 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1970:RFP

[Mol70] C. B. Moler. Reviews: First Principles of Numerical Analysis, by Burton Wendroff. American Mathematical Monthly, 77 (7):788–789, August/September 1970. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Moler:1971:MCF

[Mol71] Cleve B. Moler. Matrix computations with Fortran and paging. Technical Report STAN-CS-71-196, Computer Science Department, Stanford University, Stanford, CA, USA, 1971. 13 pp.

Moler:1972:LES

[Mol72a] C. B. Moler. Linear equation solver. Communications of the Association for Computing Machinery, 15(4):274–??, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM Algorithm 423.

Moler:1972:MCF

[Mol72b] C. B. Moler. Matrix computations with Fortran and paging. Communications of the Association for Computing Machinery, 15(4):268– 274 (or 268–270??), April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moler:1972:MGF

[Mol72c] Cleve Moler. A memory of George Forsythe. ACM SIGNUM Newsletter, 7(3):8–9, October 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1974:CRS

[Mol74] C. Moler. Cramer’s rule on 2-by-2 systems. ACM SIGNUM Newsletter, 9(4):13–14, October 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1975:DCE

[Mol75] C. B. Moler. Difficulties in computing the exponential of a matrix. In 2nd USA-Japan Computer Conference proceedings: August 26–28, 1975, Tokyo, Japan, pages 79–82 (of viii + 615). AFIPS Press, Montvale, NJ, USA, 1975.

Moler:1979:RBM

[Mol79] Cleve B. Moler. Response to Beebe by Moler [the generalized eigenvalue problem in quantum chemistry]. In Moler and Shavitt [MS78a], page 65. LCCN QD39.3.M3 W61 1978. URL

                                        http://escholarship.org/uc/
                                        item/3xb320bq. UC-32, CONF-

780878.

Moler:1980:DIM

[Mol80a] C. Moler. Design of an interactive matrix calculator. In AFIPS Conference Proceedings. 1980 National Computer Conference, pages 363–368 (of xii + 919). AFIPS Press, Montvale, NJ, USA, 1980.

Moler:1980:MIM

[Mol80b] C. B. Moler. MATLAB — an interactive matrix laboratory. Technical Report 369, Department of Mathematics and Statistics, University of New Mexico, 1980.

Moler:1980:MUG

[Mol80c] Cleve B. Moler. MATLAB user’s guide. Technical Report CS81-1, University of New Mexico. Dept. of Computer Science, November 1980. This describes use of Classic Matlab, the prototype for the very-much expanded professional Matlab from The MathWorks. Classic Matlab is no longer available.

Moler:1980:TRP

[Mol80d] Cleve B. Moler. Three research problems in numerical linear algebra. In Golub and Oliger [GO80], pages 1–18. ISBN 0-8218-0122-8, 0-8218-9237-1 (e-book). LCCN QA1 .S93 v.22.

Moler:1981:MIG

[Mol81] Cleve B. Moler. MATLAB installation guide. Technical report, University of New Mexico. Dept. of Computer Science, May 1981. This describes installation of Classic Matlab, the prototype for the very-much expanded professional Matlab from The Math-Works. Classic Matlab is no longer available.

Moler:1982:MUG

[Mol82a] C. B. Moler. MATLAB users’ guide. Technical Report CS81-1 (Revised), Department of Computer Science, University of New Mexico, 1982.

Moler:1982:DML

[Mol82b] Cleve Moler. Demonstration of a matrix laboratory. Lecture Notes in Mathematics, 909: 84–98, 1982. CODEN LN-MAA2. ISBN 3-540-11193-X (print), 3-540-38986-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/
10.1007/BFb0092962/.

Moler:1983:LNA

[Mol83] Cleve B. Moler. Learning numerical analysis and mathematics from interactive computing. In Proceedings of the 1983 American Control Conference: Sheraton-Palace Hotel, San Francisco, California, June 22, 23, 24, 1983, volume 1, pages 123–124. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September

1983. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=4788086. Three vol-

umes.

Moler:1986:MCD

[Mol86] C. B. Moler. Matrix computations on distributed memory multiprocessors. In Heath [Hea86], pages 181–195. ISBN 0-89871-215-7. LCCN QA76.5 .C61921 1986.

Moler:1988:MMV

[Mol88a] C. Moler. MATLAB: a mathematical visualization laboratory. In Digest of Papers: COMPCON Spring 88. Thirty-Third IEEE Computer Society International Conference, pages 480–481 (of xvi + 549). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0828-5. LCCN QA75.5 .C58 1988. IEEE Catalog number 88CH2539-5.

Moler:1988:SSH

[Mol88b] C. Moler. Single-user supercomputers or how I got rid of the BLAS. In Digest of Papers: COMPCON Spring 88. Thirty-Third IEEE Computer Society International Conference, pages 448–451 (of xvi + 549). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0828-5. LCCN QA75.5 .C58 1988. IEEE Catalog number 88CH2539-5.

Moler:1988:SUS

[Mol88c] Cleve B. Moler. Single-user supercomputers or how I got rid of the BLAS. In Compcon Spring ’88. Thirty-Third IEEE Computer Society International Conference, Digest of Papers, pages 448–451. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

1988. URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=4909.

Moler:1992:DGH

[Mol92] Cleve Moler. Dedication to Gene H. Golub on the occasion of his 15th birthday. SIAM Journal on Matrix Analysis and Applications, 13(1):vii–viii, January 1992. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Moler:1994:CCB

[Mol94] Cleve B. Moler. Cleve’s corner:

Benchmarks — LINPACK and MATLAB: Fame and fortune from megaflops. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Summer/Fall 1994. 3 pp.

URL http://www.mathworks.
com/company/newsletter/pdf/
sumfall94cleve.pdf.

Moler:1995:CCT

[Mol95a] Cleve B. Moler. Cleve’s corner: A tale of two numbers: With the Pentium, there is a very small chance of making a very large error. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter 1995.

                                        URL http://www.mathworks.
                                        com/company/newsletter/pdf/
                                        win95cleve.pdf.

Moler:1995:CCR

[Mol95b] Cleve B. Moler. Cleve’s corner: Random thoughts: 10^{435} years is a very long time. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Fall 1995. 2 pp.

                                        URL http://www.mathworks.
                                        com/company/newsletter/pdf/
                                        Cleve.pdf.

Moler:1995:CCQ

[Mol95c] Cleve B. Moler. Cleve’s corner: The QR algorithm: Striving for infallibility. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Summer 1995. 2 pp.

                                        URL http://www.mathworks.
                                        com/company/newsletter/pdf/
                                        sum95cleve.pdf.

Moler:1995:CCW

[Mol95d] Cleve B. Moler. Cleve’s corner: Why there isn’t a parallel MATLAB. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, 1995. URL http:

                                        //www.mathworks.com/company/
                                        newsletter/pdf/spr95cleve.
                                        pdf.

Moler:1995:TTN

[Mol95e] Cleve B. Moler. A tale of two numbers. SIAM News, 28(1):16, January 1, 1995. ISSN 0036-1437. Discusses the Intel Pentium chip divide flaw.

Moler:1996:CCF

[Mol96a] Cleve B. Moler. Cleve’s corner: Floating points: IEEE Standard unifies arithmetic model. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Fall 1996. 3 pp. URL http:/

/www.mathworks.com/company/
newsletter/pdf/Fall96Cleve.
pdf.

Moler:1996:CCG

[Mol96b] Cleve B. Moler. Cleve’s corner: Golden ODEs: New ordinary differential equation solvers for MATLAB and SIMULINK. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Summer 1996. 3 pp. URL http:

//www.mathworks.com/company/
newsletter/pdf/Sum96.Cleve.
pdf.

Moler:1997:CCW

[Mol97] Cleve B. Moler. Cleve’s corner: Are we there yet? Zero crossing and event handling for differential equations. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, 1997. 2 pp.

URL http://www.mathworks.
com/company/newsletter/pdf/
97slCleve.pdf.

Moler:1998:CCG

[Mol98a] Cleve B. Moler. Cleve’s corner: Good vibrations: eigshow helps explain eigenvalues and singular values. Technical note, The Math-Works, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter 1998. 1 pp. URL http:

                                        //www.mathworks.com/company/
                                        newsletter/clevescorner/win98cleve.
                                        shtml.

Moler:1998:CCT

[Mol98b] Cleve B. Moler. Cleve’s corner: Trigonometry is a complex subject: Revisiting inverse, complex, hyperbolic, floatingpoint trig functions. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Summer 1998. 1 pp. URL http:/

                                        /www.mathworks.com/company/
                                        newsletter/clevescorner/sum98cleve.
                                        shtml.

Moler:1999:CCOb

[Mol99a] Cleve B. Moler. Cleve’s corner: Objectively speaking: OOPS is not an apology. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter 1999. URL http:/

                                        /www.mathworks.com/company/
                                        newsletter/clevescorner/clevecorner.
                                        shtml.

Moler:1999:CCOa

[Mol99b] Cleve B. Moler. Cleve’s corner: Optimally speaking: Optimization Toolbox features new methods for large-scale problems.

Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter 1999. 1 pp. URL http:

//www.mathworks.com/company/
newsletter/clevescorner/sm99cleve.
shtml.

Moler:2000:CCF

[Mol00a] Cleve B. Moler. Cleve’s corner:

Faster finite Fourier transforms: MATLAB 6 incorporates FFTW. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter 2000. 1 pp. URL http:

//www.mathworks.com/company/
newsletter/clevescorner/winter01_
cleve.shtml.

Moler:2000:CCM

[Mol00b] Cleve B. Moler. Cleve’s corner:

MATLAB incorporates LAPACK: Increasing the speed and capabilities of matrix computation. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Winter 2000. 1 pp. URL http:

//www.mathworks.com/company/
newsletter/clevescorner/winter2000.
cleve.shtml.

Moler:2001:CCN

[Mol01] Cleve B. Moler. Cleve’s corner: Normal behavior: Ziggurat algorithm generates normally distributed random numbers. Technical note, The Math-Works, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, Spring 2001. 1 pp. URL http:

//www.mathworks.com/company/

                                        newsletter/clevescorner/spring01_
                                        cleve.shtml.

Moler:2002:CCT

[Mol02a] Cleve B. Moler. Cleve’s corner: The tetragamma function and numerical craftsmanship: MAT-LAB’s special mathematical functions rely on skills from another era. Technical note, The Math-Works, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, February 2002. 1 pp. URL http:

                                        //www.mathworks.com/company/
                                        newsletter/clevescorner/winter02_
                                        cleve.shtml.

Moler:2002:CCW

[Mol02b] Cleve B. Moler. Cleve’s corner: The world’s largest matrix computation: Google’s PageRank is an eigenvector of a matrix of order 2.7 billion. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, October 2002. 1 pp. URL http:

                                        //www.mathworks.com/company/
                                        newsletter/clevescorner/oct02_
                                        cleve.shtml.

Moler:2004:NCM

[Mol04a] Cleve B. Moler. Numerical computing with MATLAB. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2004. ISBN 0-89871-560-1. xii + 336 pp. LCCN QA297 .M625 2004.

Moler:2004:WPT

[Mol04b] Cleve B. Moler. Work in progress — teaching numerical computing with MATLAB. In ????, editor, FIE 2004: 34th Annual Conference on Frontiers in Education, 2004, pages T2H/1–T2H/2 (vol. 1). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN ???? LCCN

???? URL http://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=
&arnumber=1408512.

Moler:2006:CCP

[Mol06a] Cleve B. Moler. Cleve’s corner: Professor SVD. The Math-Works News&Notes, pages 26– 29, October 2006. URL http:

//www.mathworks.com/mason/
tag/proxy.html?dataid=8193&
fileid=35906.

Moler:2006:CCG

[Mol06b] Cleve B. Moler. Cleve’s corner:

The growth of MATLAB and The MathWorks over two decades. The MathWorks News&Notes, pages 22–24, January 2006.

Moler:2007:RSG

[Mol07] Cleve B. Moler. Recollections of a Stanford NA groupie. In Du et al. [DFM^{+}07], page ?? ISBN ???? LCCN ????

URL  http://www.stanford.
edu/group/compmath50/docs/
CompMath50Pgm21.pdf.

Moler:2008:RBV

[Mol08a] Cleve Moler. Remembering Beresford and Velvel, March

30, 2008. URL http://math.
berkeley.edu/bascd08. The Bay

Area Scientific Computing Day, BASCD08, honoring Profs. Kahan and Parlett, 29–30 March, 2008.

Moler:2008:NCM

[Mol08b] Cleve B. Moler. Numerical computing with MATLAB. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2008. ISBN 0-89871-660-8 (hardcover), 0-89871-560-1 (paperback), 0-89871-795-7 (e-book). xi + 336 pp. LCCN QA297 .M625 2008.

Moler:2011:EM

[Mol11] Cleve Moler. Experiments with MATLAB. The MathWorks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 2011.

ISBN ????  ????  pp.

                                        LCCN ???? URL http://www.
                                        mathworks.com/moler/exm/.

Moler:2013:CCG

[Mol13] Cleve B. Moler. Cleve’s corner: The Gatlinburg and Householder symposia. Technical note, The MathWorks, Inc., 3 Apple Hill Drive, Natick, MA 01760-2098, USA, ?? 2013.

??  pp.  URL https:/

                                        /www.mathworks.com/company/
                                        newsletters/articles/the-gatlinburg-
                                        and-householder-symposia.html.

Moler:2014:CCFb

[Mol14a] Cleve Moler. Cleve’s Corner: Floating point denormals, insignificant but controversial. MathWorks Web site., July 21, 2014. URL https:/

                                        /blogs.mathworks.com/cleve/

2014/07/21/floating-point-
denormals-insignificant-but-
controversial-2/.

Moler:2014:CCFa

[Mol14b] Cleve Moler. Cleve’s Corner: Floating point numbers. MathWorks Web site., July 7,

2014.  URL https://blogs.
mathworks.com/cleve/2014/07/
07/floating-point-numbers/.

Moler:2015:CCF

[Mol15a] Cleve Moler. Cleve’s Corner:

Friedrich Bauer. MathWorks Web site., June 1, 2015. URL https:

//blogs.mathworks.com/cleve/
2015/06/01/friedrich-bauer/.

Moler:2015:TPJ

[Mol15b] Cleve Moler. Technical perspective: Not just a matrix laboratory anymore. Communications of the Association for Computing Machinery, 58(10):90, October 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

cacm.acm.org/magazines/2015/
10/192374/fulltext.

Moler:2017:CCB

[Mol17a] Cleve Moler. Cleve’s corner: Bug in half-precision floating point object. Math-Works Web site., December 20,

2017.  URL https://blogs.
mathworks.com/cleve/2017/12/
20/bug-in-half-precision-floating-
point-object/. See [Mol17b].

Moler:2017:CCH

[Mol17b] Cleve Moler. Cleve’s corner: “half precision” 16-bit floating point arithmetic. MathWorks Web site., May 8, 2017. URL https:/

/blogs.mathworks.com/cleve/
2017/05/08/half-precision-
16-bit-floating-point-arithmetic/

. See bug fix [Mol17a].

Moler:2017:CCQ

[Mol17c] Cleve Moler. Cleve’s corner: Quadruple precision, 128bit floating point arithmetic. MathWorks Web site., May 22,

2017.  URL https://blogs.
mathworks.com/cleve/2017/05/
22/quadruple-precision-128-
bit-floating-point-arithmetic/

.

Moler:2018:CCV

[Mol18a] Cleve Moler. Cleve’s corner: VAXBARN restores vibrating membrane on Ardent Titan 1. MathWorks Web site., December 31, 2018. URL https:/

/blogs.mathworks.com/cleve/
2018/12/31/vaxbarn-restores-
vibrating-membrane-on-ardent-
titan/#comment-7949.

Moler:2018:HMU

[Mol18b] Cleve Moler. The historic MAT-LAB Users’ Guide. Web site, February 5, 2018. URL https:

//blogs.mathworks.com/cleve/
2018/02/05/the-historic-matlab-
users-guide/.

Moler:2019:CCC

[Mol19a] Cleve Moler. Cleve s corner: Cleve Moler on mathematics and computing. MathWorks Web site., June 25, 2019. URL

http://feeds.feedburner.com/
mathworks/moler.

Moler:2019:CCF

[Mol19b] Cleve Moler. Cleve’s corner:

Floating point arithmetic before IEEE 754. MathWorks Web site., January 18, 2019. URL https:

//blogs.mathworks.com/cleve/
2019/01/18/floating-point-
arithmetic-before-ieee-754/.

Moler:2019:CCV

[Mol19c] Cleve Moler. Cleve’s corner: Variable format half precision floating point arithmetic. MathWorks Web site., January 16, 2019. URL https:

//blogs.mathworks.com/cleve/
2019/01/16/variable-format-
half-precision-floating-point-
arithmetic/.

Moler:2019:CL

[Mol19d] Cleve Moler. Cleve s laboratory. Web site., November

8, 2019. URL https://www.
mathworks.com/matlabcentral/
fileexchange/59085-cleve_s-
laboratory.

Moler:1968:BEE

[MP68] C. B. Moler and L. E. Payne.

Bounds for eigenvalues and eigenvectors of symmetric operators. SIAM Journal on Numerical Analysis, 5(1):64–70, March 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moler:1994:NAB

[MRS94] C. B. Moler, M. W. Reichelt, and L. F. Shampine. The numerical analytic basis for MathWorks simulations tools. In J. Halin, W. Karplus, and R. Rimane, editors, CISS. First Joint Conference of International Simulation Societies Proceedings, pages 18–20 (of xvii + 816). ????, ????, 1994.

Moler:1970:ISR

[MS70a] C. B. Moler and L. P. Solomon. Integrating square roots. Communications of the Association for Computing Machinery, 13(9): 556–557, September 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moler:1970:USN

[MS70b] C. B. Moler and L. P. Solomon. Use of splines and numerical integration in geometrical acoustics. Journal of the Acoustical Society of America, 48(3):739–744, September 1970. CODEN JAS-MAN. ISSN 0001-4966.

Moler:1971:AGM

[MS71] Cleve B. Moler and Gilbert W. Stewart. An algorithm for the generalized matrix eigenvalue problem Ax = λBx. Technical Report STAN-CS-71-232, Computer Science Department, Stanford University, Stanford, CA, USA, 1971. 50 pp. Issued jointly as report CNA 32 by the Center for Numerical Analysis, the University of Texas at Austin.

Moler:1972:AGM

[MS72] C. B. Moler and G. W. Stewart. An algorithm for generalized matrix eigenvalue problems. Report TR-3, University of Michigan, Ann Arbor, MI, USA, February 1972. 50 pp.

Moler:1973:AGM

[MS73] C. B. Moler and G. W. Stewart.

An algorithm for generalized matrix eigenvalue problems. SIAM Journal on Numerical Analysis, 10(2):241–256, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moler:1978:NAC

[MS78a] Cleve B. Moler and Isaiah Shavitt, editors. Numerical algorithms in chemistry: algebraic methods: report on the workshop, August 9–11, 1978, LBL-8158. Lawrence Berkeley Laboratory, University of California, Berkeley, CA, USA, 1978. LCCN QD39.3.M3 W61 1978. URL

http://escholarship.org/uc/
item/3xb320bq. UC-32, CONF-

780878.

Moler:1978:HFA

[MS78b] Cleve B. Moler and G. W. Stewart. On the Householder–Fox algorithm for decomposing a projection. Journal of Computational Physics, 28(1):82–91, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0021999178900487.

Moler:1979:EMF

[MS79] Cleve B. Moler and G. W.

(Gilbert W.) Stewart. An efficient matrix factorization for digital image processing. Report LA-7637-MS, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, January 1979. 15 pp.

Moler:1978:DWC

[MV78a] C. Moler and C. Vanloan. 19 dubious ways to compute exponential of a matrix. SIAM Review, 20 (4):801–836, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moler:1978:NDW

[MV78b] Cleve Moler and Charles Van Loan. Nineteen dubious ways to compute the exponential of a matrix. SIAM Review, 20(4): 801–836, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moler:1979:NDW

[MV79] C. Moler and C. Van Loan. Nineteen dubious ways to compute the exponential of a matrix. In Isaac L. Auerbach, editor, The Auerbach annual, 1979, best computer papers, pages 237– 281 (of x + 334). North-Hol-land, Amsterdam, The Netherlands, 1979. ISBN 0-444-00350-9. LCCN QA76.24.A8 1979.

Moler:2003:NDW

[MV03] Cleve Moler and Charles Van Loan. Nineteen dubious ways to compute the exponential of a matrix, twenty-five years later. SIAM Review, 45(1):3–49, March 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/
article/41801.

Nickel:1967:ERP

[Nic67] Karl L. E. Nickel. Extension of a recent paper by Fox, Henrici and Moler on eigenvalues of elliptic operators. SIAM Journal on Numerical Analysis, 4(4): 483–488, December 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (elec-

tronic).  URL https://www.
jstor.org/stable/2949713.

Petzold:2001:GEC

[PAB^{+}01] L. R. Petzold, U. Ascher, H. T.

Banks, J. Crowley, W. Gander, L. Greengard, M. Heath, A. Lumsdaine, C. Moler, T. Oden, R. Schnabel, K. Stewart, and A. Trefethen. Graduate education in computational science and engineering. SIAM Review, 43(1): 163–177, ???? 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Phillips:1989:DGW

[PPM^{+}89] D. Phillips, M. Pique, C. Moler, J. Torborg, and D. Greenberg. Distributed graphics: where to draw the lines? ACM SIG-GRAPH—Computer Graphics, 23(5):257–280, December 1989. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Smith:1976:MER

[SBD^{+}76] B. T. Smith, J. M. Boyle, J. J. Dongarra, B. S. Garbow, Y. Ikebe, V. C. Klema, and C. B. Moler. Matrix Eigensystem Routines—EISPACK Guide, volume 6 of Lecture Notes in Computer Science, Editors: G. Goos and J. Hartmanis. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1976. ISBN 0-387-06710-8, 3-540-07546-1. vii + 387 pp. LCCN QA193 .M37, QA267.A1 L43 no.6.

Strang:2022:LCE

[SM22] Gilbert Strang and Cleve Moler. LU and CR elimination. SIAM Review, 64(1):181–190, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https://

                                        epubs.siam.org/doi/10.1137/
                                        20M1358694.

Stenger:1968:TC

[Ste68] Frank Stenger. A trap in computations. ACM SIGNUM Newsletter, 3(3):2, July 1968. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [Mol69a].

USENIX:1994:PUSb

[USE94] USENIX Association, editor. Proceedings of the USENIX Symposium on Very High Level Languages (VHLL): October 26–28, 1994, Santa Fe, New Mexico, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.